### v2

# Toward Conversational Hungarian Speech Recognition: Introducing the BEA-Large and BEA-Dialogue Datasets

Abstract
The advancement of automatic speech recognition (ASR) has been largely enhanced by extensive datasets in high-resource languages, while languages such as Hungarian remain underrepresented due to limited spontaneous and conversational corpora. To address this gap, we introduce two new datasets – BEA-Large and BEA-Dialogue – constructed from the previously unprocessed portions of the Hungarian speech corpus named BEA. BEA-Large extends BEA-Base with 255 hours of spontaneous speech from 433 speakers, enriched with detailed segment-level metadata. BEA-Dialogue, comprising 85 hours of spontaneous conversations, is a Hungarian speech corpus featuring natural dialogues partitioned into speaker-independent subsets, supporting research in conversational ASR and speaker diarization. We establish reproducible baselines on these datasets using publicly available ASR models, with the fine-tuned Fast Conformer model achieving word error rates as low as 14.18% on spontaneous and 4.8% on repeated speech. Diarization experiments yield diarization error rates between 12.46% and 17.40%, providing reference points for future improvements. The results highlight the persistent difficulty of conversational ASR, particularly due to disfluencies, overlaps, and informal speech patterns. By releasing these datasets and baselines, we aim to advance Hungarian speech technology and offer a methodological framework for developing spontaneous and conversational benchmarks in other languages.
Keywords: speech database, automatic speech recognition, spontaneous speech, evaluation
[…]
Toward Conversational Hungarian Speech Recognition: Introducing the BEA-Large and BEA-Dialogue Datasets
Máté Gedeon∗,†, Piroska Zsófia Barta∗,†, Péter Mihajlik∗,‡, Tekla Etelka Gráczi, Anna Kohári, Katalin Mády
Dept. of Telecommunications and Artificial Intelligence, Budapest University of Technology and Economics, Hungary ELTE Research Centre for Linguistics, Hungary Speechtex Ltd. {gedeonm, piroskazsofia.barta}@edu.bme.hu
{mihajlik.peter, graczi.tekla.etelka, kohari.anna, mady}@nytud.hu
Abstract content
1.  Introduction
The field of automatic speech recognition (ASR) has been fundamentally transformed by the availability of large-scale training datasets Panayotov et al. (2015); Pratap et al. (2020). However, this revolution has predominantly benefited well-resourced languages, leaving low-resource languages like Hungarian significantly underrepresented in the modern ASR ecosystem Besacier et al. (2014). While recent advances in self-supervised learning and multilingual models have shown promise for low-resource scenarios, the fundamental challenge remains: the scarcity of high-quality, diverse speech corpora that capture the full spectrum of natural human communication Mihajlik et al. (2024).
Hungarian, with its morphologically rich and agglutinative nature, poses a significant challenge even for large multilingual models. To improve the results of these models via fine-tuning, or by training our own, having high-quality spontaneous speech as well as conversational datasets is vital Mihajlik et al. (2010). BEA Neuberger et al. (2014) is a large Hungarian speech database with the aim, among others, to provide material for research purposes in various fields. The complete 300-hour dataset consists of eight types of speech sessions, including repeated and spontaneous speech, from almost 500 speakers of varying age, gender, and educational background.
Although the BEA Base dataset Mihajlik et al. (2022), comprising 140 speakers, has been released as a benchmark for ASR of spontaneous Hungarian, a considerable part of the BEA dataset has not yet been processed and subjected to benchmarking.
In this work, we leverage the remaining BEA recordings to create two new datasets and provide comprehensive baselines for Hungarian conversational ASR. The first dataset, BEA-Large, extends and refines BEA-Base by including additional training data and more extensive segment-level metadata. BEA-Large thus offers an extended corpus of spontaneous Hungarian, with fine-grained annotations to support research in ASR and related areas. The second dataset, BEA-Dialogue, mainly contains dialogues from BEA-Large. Alongside the aforementioned datasets, we provide baseline ASR results for both, using publicly available models to ensure reproducibility. These resources not only advance Hungarian speech technology but also provide a methodological framework for similar efforts in other low-resource languages. By releasing these datasets along with comprehensive baselines, we aim to catalyze research in spontaneous Hungarian speech recognition while contributing to the broader goal of democratizing speech technology across linguistic boundaries.
Our primary contributions are:
  • An extended dataset (BEA-Large) comprising 255 hours of speech from 433 speakers, substantially expanding the available Hungarian spontaneous speech training data with enriched metadata
  • A specialized conversational speech dataset (BEA-Dialogue) featuring 85 hours of natural dialogues, addressing the critical shortage of Hungarian dialogue data for conversational ASR and speaker diarization research
  • Systematic benchmarking using publicly available ASR models to ensure reproducibility and establish performance baselines
The structure of the paper is as follows. Section 2 reviews related work. Sections 3 and 4 describe the construction of the two datasets, followed by ASR baseline results in Sections 5 and 6, and diarization results on BEA-Dialogue in Section 7. Finally, Section 8 summarizes the findings and concludes the paper.
2.  Related work
Concerning the international situation of spontaneous speech datasets, there are several resources for a range of languages and domains. However, a significant portion of these corpora remains unavailable for public use, especially for low-resource languages. Among the publicly accessible English datasets, SSSD Sheikh et al. (2025) is designed for dialogue research and features 727 hours of spontaneous English conversations between randomly matched speaker pairs. CASPER Xiao et al. (2025) provides over 100 hours of unscripted English dialogues.
Beyond English, the GRASS corpus Schuppler et al. (2014)—the only existing resource for Austrian German for years Linke et al. (2022)—which contains approximately 19 hours of dyadic speech; and ES-Port García-Sardiña et al. (2018), comprising 40 hours of Spanish dialogues from technical support calls. The RAMC corpus Yang et al. (2022) encompasses 180 hours of conversational speech recorded via telephone channels. The Verbmobil dialogue corpus Weilhammer et al. (2002) covers multilingual data in German, English, and Japanese. The Kiel Corpus of Spoken German Kohler et al. (2017) includes over 8 hours of spontaneous speech from 64 speakers.
For specialized domains, RescueSpeech Sagar et al. (2023) focuses on the search and rescue context, containing approximately 2 hours of annotated German speech from simulated exercises. The SXUCorpus Herms et al. (2016) offers spontaneous speech components in the Upper Saxon German dialect. The Portuguese CORAA corpus Junior et al. (2021) comprises at least 189 hours of spontaneous speech from conversations, monologues, dialogues, and interviews. The CSJ corpus Uchimoto et al. (2007), while primarily consisting of monologues, also features spontaneous speech. The Norwegian Parliamentary Speech Corpus Solberg and Ortiz (2022) includes unscripted parliamentary sessions, while the Korean Corpus of Spontaneous Speech Yun et al. (2015) contains recordings from interviews with 40 speakers.
As discussed by Mihajlik et al. (2024), despite the relatively large cumulative volume of Hungarian speech data, inconsistencies in annotation practices and restricted accessibility hinder their simultaneous use—both for general-purpose ASR and for specialized tasks such as conversational ASR. Among the monolingual Hungarian resources, BUSZI ($\sim$600h, 250 spk; Kontra et al. (1997)) and SzöSzi (370h, 163 spk; Kontra et al. (2016)) stand out as extensive sociolinguistic interview collections – unfortunately, their access is strongly limited. HuComTech ($\sim$50h, 112 spk; Pápay et al. (2011)) provides valuable audiovisual recordings for multimodal communication studies, while BEKK (20h, 56 spk; Bodó et al. (2017)) contributes spontaneous, student-recorded conversations.
3.  Construction of BEA-Large
The BEA-Large dataset builds upon the original BEA-Base corpus by extending its training portion with an additional subset, BEA-Extension (later referred to as train-293), with each segment corresponding to a single utterance of a given speaker. The dev and eval partitions remain identical to those in BEA-Base to preserve consistency and comparability across experiments.
train-114 train-293
# speakers [f|m] 69 | 45 184 | 109
# segments 69,176 196,981
# words 555,322 1,622,151
# chars 3,310,493 9,550,276
duration [h] 67.95 177.4
Table 1: Metadata for the disjoint training sets of BEA-Large.
The additional training set introduces new metadata attributes that were absent or incomplete in BEA-Base, including speaker age, gender, occupation, and the module (e.g., interview, discourse) from which each segment originates. In addition to the unique identifier of the target speaker associated with each recording, the dataset now explicitly labels each speaker’s role according to the BEA-Base taxonomy: SPK (target speaker), EXP (experiment leader), or DP (discourse partner).
The train-293 subset was compiled from recordings of 293 target speakers who do not appear in BEA-Base, ensuring that both training sets can be seamlessly combined without overlap. As summarized in Table 2, data were drawn from multiple modules detailed by Mihajlik et al. (2022), with the repeat and readsent modules excluded to maintain structural alignment with the train-114 set. Acoustic preprocessing followed the same protocol as in BEA-Base.
Module SPK [%] EXP [%] DP [%]
repeat
readsent
interview 16.45 3.53
opinion 12.86 5.56
summhist 3.80 0.36
summplant 3.26 0.46
discourse 21.73 18.02 8.39
readtext 5.46 0.14
Table 2: Composition of the train-293 set (in percentage of total duration).
When combined with train-114, the resulting BEA-Large training corpus contains approximately three times as much data as BEA-Base. The duration of the segments in BEA-Large is shown in Figure 2, with approximately 99.6% of the segments in the database have a duration of less than 20 seconds. Figure 1 illustrates the age distribution of speakers in the BEA-Large training set compared to that of train-114.
Figure 1: Comparison of the age distributions in the train sets of BEA-Large and Base.
Figure 2: Duration of the segments in the train sets of BEA-Large and Base.
Model Dataset dev-repet eval-repet dev-spont eval-spont
WER CER WER CER WER CER WER CER
whisper-large-v3 zero-shot 13.31 2.74 13.73 2.83 20.79 9.22 21.28 9.31
whisper-large-v2 zero-shot 18.56 4.47 21.26 4.46 34.71 19.67 32.56 18.61
whisper-medium zero-shot 22.7 5.27 23.84 6.10 37.23 20.13 39.19 21.71
f_conformer_ctc train-114 7.01 1.54 7.05 1.82 18.26 6.04 19.22 6.37
train-293 14.60 2.83 14.98 3.30 15.60 5.18 16.33 5.34
train-114 + train-293 5.21 1.16 4.80 1.20 13.52 4.41 14.18 4.56
Table 3: WER and CER [%] for zero-shot and fine-tuned models on BEA-Large under different configurations.
4.  Construction of BEA-Dialogue
A dedicated dialogue corpus, BEA-Dialogue, was derived from the recordings of 242 speakers that had not previously been included in BEA-Base.
In the original dataset, utterances were organized by the target speaker, with each module stored in a TextGrid file containing time-aligned transcriptions for all speakers appearing in that module. To create BEA-Dialogue, utterances were first extracted along with their timestamps and speaker labels (SPK, EXP, DP).
Candidate cut points for shorter dialogue segments were then identified by detecting silent intervals—regions that did not overlap with any utterance, except possibly at their boundaries. Using these silence-based boundaries, utterances from multiple speakers were grouped into coherent dialogue units according to the silent regions separating them. These smaller units were subsequently merged into larger dialogue segments, each with a target duration of 30 seconds. Figure 3 shows the histogram of the durations after the procedure.
Figure 3: Duration of the segments in BEA-Dialogue.
Beyond the target speakers, the dataset includes the voices of five additional female speakers (fem1–fem5) and two male speakers (male1–male2), who served as experiment leaders and discourse partners. To construct BEA-Dialogue, which provides fully disjoint training, development, and evaluation sets across all speaker roles, the data were partitioned based on three experiment leaders: fem1, fem3, and fem4.
Since these experiment leaders also acted as discourse partners for different target speakers, we excluded dialogue segments from the discourse module in which the target speaker’s discourse partner was one of the three experiment leaders. The combinations of experiment leaders and discourse partners across all available data are shown in Table 4, along with the corresponding number of dialogue segments belonging to each pair.
EXP
fem1 fem3 fem4
DP fem1 286 852
fem2 1751 122
fem3 4415 546
fem4 4418
fem5 1662 367 816
male1 983 88 588
male2 46
Table 4: Number of discourse segments by experiment leader (EXP) and discourse partner (DP) across all dialogue segments.
The training set includes dialogue segments where the experiment leader is fem1 and the discourse partner is either fem2 or male2, as well as all segments from other modules featuring fem1. The development and evaluation sets contain dialogues where the experiment leaders are fem3 and fem4, with discourse partners fem5 and male1, respectively. Table 5 summarizes the metadata of BEA-Dialogue, which represents the largest set that can be derived from the data without violating speaker independence across subsets with respect to the three speaker roles (SPK, EXP, DP).
BEA-Dialogue
Train Dev Eval
# Speakers [f|m] 121 | 67 3 | 6 29 | 16
# Segments 9,179 577 1,906
# Words 532,732 34,056 105,472
# Characters 3,217,617 206,740 641,628
Avg. # Speakers / Seg 1.77 1.92 1.61
Avg. # Utterances / Seg 10.99 8.68 9.74
Avg. Seg Duration [s] 26.23 26.31 26.09
Overlap Duration [h] 3.28 0.29 0.40
Total Duration [h] 66.87 4.22 13.81
Table 5: Metadata of the BEA-Dialogue dataset.
5.  Baseline ASR results for BEA-Large
To establish reproducible baselines, we fine-tuned a potent yet relatively small (120M parameters) model publicly available in the NeMo toolkit Kuchaiev et al. (2019)STT En Fast Conformer-CTC Large  [^1] https://huggingface.co/nvidia/stt_en_fastconformer_ctc_large (f_conformer_ctc) – on three different training sets: train-114 (BEA-Base), train-293 (BEA-Extension) and train-114 + train-293 (BEA-Large).
Performance was evaluated using two standard metrics: Word Error Rate (WER) and Character Error Rate (CER). The results are summarized in Table 3. Zero-shot benchmarks obtained using the Speechbrain toolkit Ravanelli et al. (2021, 2024) with three Whisper models [^2] https://github.com/openai/whisperwhisper-large-v3, whisper-large-v2, and whisper-medium Radford et al. (2022) — are also presented in Table 3.
The Fast Conformer model fine-tuned on less than 250 hours of Hungarian speech consistently outperformed all Whisper models, across both the repetitive and spontaneous subsets of BEA-Large.
In both training partitions of BEA-Large, the voices of the seven individuals acting as experiment leaders and discourse partners are notably overrepresented. To obtain two speaker-independent subsets, we extracted utterances from the two training sets belonging exclusively to the target speakers. We constructed three setups from the two speaker-independent supplementary training sets: using each dataset individually, one at a time, and using both datasets merged. Notably, the WER and CER increased when the training sets were limited to the utterances of the target speakers.
Using the new train-293 as the only training set resulted in higher WER and CER on dev-repet and eval-repet compared to using train-114. This increase may be caused by data leakage between the original BEA-Base splits, which will be the subject of further investigation.
The best results were achieved by using the two full components – train-114 and train-293 – of the database yielding a WER of 14.18% and a CER of 4.56% on the most important spontaneous evaluation subset, indicating that using the voice of the experiment leaders and discourse partners does not have a negative impact on the evaluation results obtained on a speaker-independent evaluation set.
6.  Baseline ASR results for BEA-Dialogue
For the BEA-Dialogue dataset, we trained a similar set of models to those used for BEA-Large, complemented with fine-tuned Fast Conformer models. During training, we employed Serialized Output Training (SOT) Kanda et al. (2020), marking speaker transitions explicitly with a <sc> (speaker change) token. These tokens were inserted to reflect the structure of an ideal dialogue transcript. For example (English translations in parentheses):
Hogy vagy? <sc> Köszönöm, jól. És te? <sc> Én is, köszönöm. <sc> Örülök neki. (How are you? <sc> I’m good, thanks. And you? <sc> Me too, thanks. <sc> I’m glad.)
The utterance of each speaker is maintained as an uninterrupted sequence, even when there is overlap in speech. This allows the model to learn to recognize speaker boundaries while preserving the linguistic integrity of each turn.
Evaluation followed the same criteria as in BEA-Large: WER and CER, complemented by the concatenated minimum-permutation WER (cpWER) and cpCER, which minimize errors across all possible permutations of speakers separated by <sc> tokens. Since some segments contained more than ten speaker changes, exhaustive permutation computation was infeasible. To address this, we adopted a hybrid strategy: computing all permutations for up to seven speaker changes and applying beam search to achieve near-optimal results for higher counts. For fine-tuned models, we additionally report the speaker change accuracy (scAcc)—the proportion of utterances in which the predicted number of speaker transitions matches the reference (i.e., occurrences of <sc> tokens).
Table 6 summarizes the obtained results.
Model dev eval
WER cpWER CER cpCER scAcc WER cpWER CER cpCER scAcc
whisper-large-v3 21.19 21.04 12.74 12.56 22.21 22.13 12.27 12.18
whisper-large-v2 19.65 19.42 9.84 9.58 24.50 24.42 13.13 13.05
whisper-medium 25.45 25.27 12.61 12.42 29.21 29.12 14.71 14.63
f_conformer_ctc 19.69 19.53 7.95 7.78 69.32 20.56 20.44 9.11 9.00 82.16
Table 6: ASR accuracy (%) comparison on BEA-Dialogue.
As shown in Table 6, the fine-tuned f_conformer_ctc model achieves lower error rates on the eval set, demonstrating the benefits of domain-specific adaptation. Interestingly, however, on the dev set, the Whisper-large-v2 model outperforms both its v3 variant and the fine-tuned model. Nevertheless, the fact that an English model fine-tuned on fewer than 70 hours of Hungarian data can surpass state-of-the-art systems underscores the importance of corpora such as BEA-Dialogue for advancing Hungarian conversational ASR. These findings highlight the current limitations of state-of-the-art models when applied to spontaneous, real-world speech.
The ratio of character error rate (CER) to word error rate (WER) is notably higher than typically observed in non-conversational datasets, suggesting that fine-grained, character-level deviations occur more frequently in spontaneous dialogue.
Figure 4 shows the distribution of the number of speaker turns per segment for each split. As shown, most segments contain no or only one speaker change, making the task largely comparable to single-speaker spontaneous speech recognition. However, the ones with a higher number of turns substantially increase task complexity.
Figure 4: Distribution of speaker changes per segment in BEA-Dialogue.
7.  Baseline diarization results for BEA-Dialogue
We establish baseline diarization performance for the BEA-Dialogue dataset using its dev and eval splits. The goal of these baselines is to provide reference performance levels for future work on speaker segmentation.
Two state-of-the-art diarization systems were evaluated: pyannote.audio Bredin (2023) and Sortformer Park et al. (2024). Both models were used in their pre-trained configurations without additional fine-tuning. The pyannote.audio model represents a hybrid neural approach that combines segmentation and clustering stages, while Sortformer adopts a transformer-based end-to-end architecture with a sorting-based loss function to better handle overlapping speech. The exact pre-trained checkpoints used are: speaker-diarization-3.1333https://huggingface.co/pyannote/speaker-diarization-3.1 and diar_sortformer_4spk-v1444https://huggingface.co/nvidia/diar_sortformer_4spk-v1.
Table 7 reports the average Diarization Error Rate (DER) obtained by each model on the two dataset subsets.
Model DER (%)
dev eval
Pyannote 14.09 17.40
Sortformer 12.46 15.11
Table 7: Average DER comparison on BEA-Dialogue.
Figure 5 illustrates the distribution of DER values across individual dialogue segments in the eval subset. Both systems exhibit similar overall trends.
Figure 5: DER distribution on BEA-Dialogue (eval).
8.  Conclusion
This work addresses the shortage of high-quality spontaneous and conversational speech data for Hungarian by introducing two substantial resources: BEA-Large and BEA-Dialogue. BEA-Large extends the original BEA-Base corpus with approximately 178 hours of additional speech from 293 speakers, enriched with segment-level metadata such as age, gender, occupation, and speaker role. This expansion nearly triples the amount of available training data while maintaining compatibility with the original benchmark splits.
BEA-Dialogue, comprising 85 hours of natural conversations, represents one of the largest Hungarian datasets explicitly designed for conversational ASR and speaker diarization research. The dataset includes predefined splits that ensure complete speaker independence across all conversational roles.
Our baseline experiments yield several key findings. Fine-tuned Fast Conformer models achieved WERs as low as 14.18% on spontaneous speech and 4.8% on repeated speech when trained on the combined datasets, highlighting the benefits of including data from both experiment leaders and discourse partners. In conversational ASR, the relatively high CER-to-WER ratio compared to monologue speech suggests that spontaneous dialogue introduces additional challenges—arising from phenomena such as overlapping speech, rapid turn-taking, and interaction-driven disfluency patterns that are not present in single-speaker settings. The speaker diarization baselines, with DERs ranging from 12.46% to 17.40%, establish solid reference points for future advancements.
By releasing these datasets together with reproducible baselines built on publicly available models, we aim to accelerate research in Hungarian speech technology and provide a methodological blueprint for similar initiatives in other low-resource languages. Future work will focus on advanced methods for overlap handling, speaker change detection, and context-aware modeling, with the goal of improving both ASR and diarization performance in conversational scenarios.
Acknowledgment
Project No. 2025-2.1.2-EKÖP-KDP-2025-00005 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the EKÖP_KDP-25-1-BME-21 funding scheme.
The work was also partially supported by the Hungarian NRDI Fund through the projects NKFIH K143075 and K135038, NKFIH-828- 2/2021(MILAB).
9.  Bibliographical References
  • L. Besacier, E. Barnard, A. Karpov, and T. Schultz (2014) Automatic speech recognition for under-resourced languages: a survey. Speech Communication 56, pp. 85–100. External Links: ISSN 0167-6393, Document, Link Cited by: §1.
  • H. Bredin (2023) Pyannote.audio 2.1 speaker diarization pipeline: principle, benchmark, and recipe. In Interspeech 2023, pp. 1983–1987. External Links: Document, ISSN 2958-1796 Cited by: §7.
  • N. Kanda, Y. Gaur, X. Wang, Z. Meng, and T. Yoshioka (2020) Serialized output training for end-to-end overlapped speech recognition. In Interspeech 2020, pp. 2797–2801. External Links: Document, ISSN 2958-1796 Cited by: §6.
  • O. Kuchaiev, J. Li, H. Nguyen, O. Hrinchuk, R. Leary, B. Ginsburg, S. Kriman, S. Beliaev, V. Lavrukhin, J. Cook, P. Castonguay, M. Popova, J. Huang, and J. M. Cohen (2019) NeMo: a toolkit for building ai applications using neural modules. External Links: 1909.09577, Link Cited by: §5.
  • J. Linke, P. N. Garner, G. Kubin, and B. Schuppler (2022) Conversational speech recognition needs data? experiments with Austrian German. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, N. Calzolari, F. Béchet, P. Blache, K. Choukri, C. Cieri, T. Declerck, S. Goggi, H. Isahara, B. Maegaard, J. Mariani, H. Mazo, J. Odijk, and S. Piperidis (Eds.), Marseille, France, pp. 4684–4691. External Links: Link Cited by: §2.
  • P. Mihajlik, K. Mády, A. Kohári, F. S. Fruzsina, G. Kiss, T. E. Gráczi, and A. S. Doğruöz (2024) Is spoken Hungarian low-resource?: a quantitative survey of Hungarian speech data sets. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), N. Calzolari, M. Kan, V. Hoste, A. Lenci, S. Sakti, and N. Xue (Eds.), Torino, Italia, pp. 9382–9388. External Links: Link Cited by: §1, §2.
  • P. Mihajlik, Z. Tuske, B. Tarján, B. Németh, and T. Fegyó (2010) Improved recognition of spontaneous hungarian speech—morphological and acoustic modeling techniques for a less resourced task. IEEE Transactions on Audio, Speech, and Language Processing 18 (6), pp. 1588–1600. External Links: Document Cited by: §1.
  • T. J. Park, I. Medennikov, K. Dhawan, W. Wang, H. Huang, N. R. Koluguri, K. C. Puvvada, J. Balam, and B. Ginsburg (2024) Sortformer: a novel approach for permutation-resolved speaker supervision in speech-to-text systems. External Links: Link Cited by: §7.
  • A. Radford, J. W. Kim, T. Xu, G. Brockman, C. McLeavey, and I. Sutskever (2022) Robust speech recognition via large-scale weak supervision. External Links: 2212.04356, Link Cited by: §5.
  • M. Ravanelli, T. Parcollet, A. Moumen, S. de Langen, C. Subakan, P. Plantinga, Y. Wang, P. Mousavi, L. D. Libera, A. Ploujnikov, F. Paissan, D. Borra, S. Zaiem, Z. Zhao, S. Zhang, G. Karakasidis, S. Yeh, P. Champion, A. Rouhe, R. Braun, F. Mai, J. Zuluaga-Gomez, S. M. Mousavi, A. Nautsch, H. Nguyen, X. Liu, S. Sagar, J. Duret, S. Mdhaffar, G. Laperrière, M. Rouvier, R. D. Mori, and Y. Estève (2024) Open-source conversational ai with speechbrain 1.0. Journal of Machine Learning Research 25 (333). External Links: Link Cited by: §5.
  • M. Ravanelli, T. Parcollet, P. Plantinga, A. Rouhe, S. Cornell, L. Lugosch, C. Subakan, N. Dawalatabad, A. Heba, J. Zhong, J. Chou, S. Yeh, S. Fu, C. Liao, E. Rastorgueva, F. Grondin, W. Aris, H. Na, Y. Gao, R. D. Mori, and Y. Bengio (2021) SpeechBrain: a general-purpose speech toolkit. Note: arXiv:2106.04624 External Links: 2106.04624 Cited by: §5.
10.  Language Resource References
[…]